Pi_{}eciogodzinówka, dzień trzeci

luty 2005

7. Niech p > 3 b_{}edzie liczb _{}a pierwsz _{}a zaś m = \frac{4^{p}−1}{3} . Wykazać, że 2^{m−1} ≡ 1 mod m.

8. Punkty A, P, Q, R i S leż _{}a na okr_{}egu w tej właśnie kolejności, przy czym

∡P AQ = ∡QAR = ∡RAS.

Wykazać, że

AR(AP + AR) = AQ(AQ + AS).

9. Wykazać, że dla x, y, z ∈ ℝ^{+} zachodzi

\frac{xyz}{x^{3} + y^{3} + xyz} + \frac{xyz}{y^{3} + z^{3} + xyz} + \frac{xyz}{z^{3} + x^{3} + xyz} 1.